OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (ujjwal):

add both equations. y will be eliminated. you will get value of x.

OpenStudy (anonymous):

then i get 2(16 +y) +y = 2 ?

OpenStudy (anonymous):

x - y = 16 X-Y=16 2x + y = 2 6-Y=16 3X=18 -6 -6 3X/3 =18/3 X=6 -Y=10 Y=-10

OpenStudy (ujjwal):

After adding, you don't get that. y will be eliminated and you will get 3x=18 @AshhSmith

OpenStudy (anonymous):

this is substitution method ?

OpenStudy (ajprincess):

no it is elimination method. u wanna do t using substitution method?

OpenStudy (anonymous):

no this fine if it get the same answer i just wasn't sure .

OpenStudy (ajprincess):

u will get the same answer in both the methods.